Simplify
(20 x 10 -4 power)(4 x 10 ninth power)
______________________________________
3,200 x 10 -3power

Answer by stanbon(75887)

Simplify
(20 x 10 -4 power)(4 x 10 ninth power)
______________________________________
3,200 x 10 -3power
============
(20E-4)(4E10)/(3200E-3) = 80E6/3200E-3 = 1E6/40E-3 = 1E9/40
= (1E9)/(4E1
= 0.25E8
= 2.5E7
============
Cheers,
Stan H.